a parallel plate capacitor is being charged by battery.Show that the displacement current across an area in the region between the plates and parallel to it is equal to the conduction current in the connecting wires.

Grade:12

2 Answers

Subodh Wasnik
4 years ago
Charge on plates
Ic=dq/dt
 
Displacement current through plates is given by
Id=ϵo dΦE/dt
 
Using Gaussian theorem,
ΦE=q/ϵo
 
Id=ϵo d(q/ϵo)/dt
Id=dq/dt
 
Therefore,both displacement currentand conduction curent are equal
Vikas TU
4 years ago
Dear student 
Good Luck